OpenStudy (anonymous):

really need help. please help. Which of the following is the solution to the equation 8^(y + 2) = 2^y ?

OpenStudy (unklerhaukus):

if you can get both sides of the equation to become 2^(something) = 2^(something else) you can simplify something = somethingelse

OpenStudy (anonymous):

not ringing any bells cause its so early in the morning......

OpenStudy (unklerhaukus):

can you express 8 as a power of two?

OpenStudy (unklerhaukus):

have you had your breakfast yet?

OpenStudy (austinl):

Breakfast is an important part of mathematics, vital even. You should have your Wheaties before trying to do algebra! But.... \[2^x=8\] x=?

OpenStudy (unklerhaukus):

\[2^x=\overbrace{2\times2\times\dots\times2}^{x\text{ times}}\]